OpenStudy (lollylau):

Differentiate (3y/x^2)^x, no using computers to help.

OpenStudy (lollylau):

time limit: 5 mins, starting now.

OpenStudy (lollylau):

time limit extended 5 mins more.

OpenStudy (anonymous):

a quiz for us?

OpenStudy (zarkon):

you should probably specify which variable you want to differentiate with respect to.

OpenStudy (lollylau):

X

OpenStudy (akshay_budhkar):

5 minutes extra now :P lol

OpenStudy (lollylau):

I'm dying.

OpenStudy (zarkon):

sorry to hear that

OpenStudy (lollylau):

Imagine not using a computer when differentiating.

OpenStudy (turingtest):

I got my answer because it's similar to the last. I'm not sure it's right, but I wouldn't want to kill your fun if you think I'm cheating.

OpenStudy (lollylau):

you can type it down

OpenStudy (lollylau):

...

OpenStudy (turingtest):

I assume that is a yes...

OpenStudy (lollylau):

How did you start?

OpenStudy (lollylau):

ln y = x ln (3y/x^2)

OpenStudy (lollylau):

y'/y=a lot of stuff

OpenStudy (turingtest):

I didn't take the log, but that would've helped. actually now I'm getting\[y=\sqrt{\frac{3y}{e^3}}\]do you know the other is right?

OpenStudy (lollylau):

no wait

OpenStudy (turingtest):

that should be an x under the radical

OpenStudy (lollylau):

Lemme type the question as clearly as I could, since I didn't

OpenStudy (lollylau):

y = (3y/x^2)^x, differentiate y.

OpenStudy (lollylau):

Im getting confused.

OpenStudy (lollylau):

I'm feeling that I'm getting there.

OpenStudy (turingtest):

Me too, I keep getting different answers. I tried to ignore all the parts that can't be zero to cut down on work, but I think I made a mistake.

OpenStudy (lollylau):

You tried to ignore the parts that can't be zero? ?

OpenStudy (zarkon):

looks like you need to use implicit differentiation

myininaya (myininaya):

\[y=(\frac{3y}{x^2})^x\] Take natural log of both sides note:also remember derivative of y is y' \[\ln(y)=x \ln(\frac{3y}{x^2})\] \[\ln(y)=x [\ln(3y)-\ln(x^2)]\] \[\ln(y)=x[\ln(3)+\ln(y)-2 \ln(x)]\] \[\frac{y'}{y}=x[0+\frac{y'}{y}-2 \cdot \frac{1}{x}]+1[\ln(3)+\ln(y)-2 \ln(x)]\] \[\frac{y'}{y}=x \cdot \frac{y'}{y}-2+\ln(3)+\ln(y)-2 \ln(x)\]

OpenStudy (lollylau):

I am.

myininaya (myininaya):

\[\frac{y'}{y}-\frac{x y'}{y}=-2+\ln(3)+\ln(y)-2 \ln(x)\] \[y'(\frac{1-x}{y})=-2 +\ln(3)+\ln(y)-2 \ln(x)\]

myininaya (myininaya):

\[y'=\frac{y}{1-x}(-2+\ln(3)+\ln(y)-2\ln(x))\]

myininaya (myininaya):

now if its multiple choice I will leave the dressing up to you

myininaya (myininaya):

dressing up as in simplifying

OpenStudy (lollylau):

y' = y(ln(3y/x^2)-2)?

OpenStudy (turingtest):

that give the same answer, but is not what I have...

OpenStudy (turingtest):

mine should have yielded what I wrote the second time, I am confused.

OpenStudy (lollylau):

y'/y = d/dx (x) ln(3y/x^2) + d/dx (ln(3y/x^2)) (x)